For the reaction

N2(g) + 3H2(g) ⇌ 2NH3(g)


Kc is 4.38 at 365 C. Calculate KP for the reaction at this temperature.
Answer #1

T= 365 oC

= (365+273) K

= 638 K

Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant

Δ n = -2

use:

Kp= Kc (RT)^Δ n

Kp = 4.38*(0.08206*638.0)^(-2)

Kp = 1.598*10^-3

Answer: 1.60*10^-3